OpenStudy (anonymous):

simultaneous equations 5p+3q=25 3p+3q=21

OpenStudy (jdoe0001):

you mean a "system of equations" :/

OpenStudy (anonymous):

umm no simultaneous equations where i think u substitute one sum into other or sthin!!

OpenStudy (jdoe0001):

hehe, ok, well, so, if you multiply the bottom one by -1, what would that give you?

OpenStudy (anonymous):

WAT!?!?! now remember u gotta find p and q

terenzreignz (terenzreignz):

Yes, @jdoe0001 is leading (very well) into that :D

OpenStudy (anonymous):

sorry ;)) i was meant to rite now i remember

terenzreignz (terenzreignz):

He is using a method called elimination, of course, you can always use substitution if that's what you're comfortable with~

OpenStudy (jdoe0001):

5p+3q=25 3p+3q=21 x (-1) => 5p+3q=25 ???+???=???